Baby Rachel wants to arrange 9 blocks in a row. How many different arrangements can she make?
1
Expert's answer
2020-12-20T16:51:08-0500

For the first place, Rachel can put any of the 9 blocks, so we have 9 options. For the second place, there will be 8 blocks left, so 8 options in this case. For the third place there will be 7 options and so on. The number of different arrangements is

"N = 9 \times 8 \times 7 \times 6 \times5 \times4 \times 3 \times 2 \times 1 = 362\,880"